Prev Next

SW&IT Meeting Minutes 2021-08-23

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20210809
LR20210809 - Bec, Matthieu 80.0
LR20210809 - Edmeades, David 80.0
LR20210809 - Grenz, Paul 40.0
LR20210809 - Hooper, Stephen 80.0
LR20210809 - Zhang, Yang 80.0
LR20210809 - Zhang, Xianyu 91.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management, ATT discussion with WRA
    • on-boarding Glenn, Brandon
    • update jenkins release, f34 build fixes
    • dms refmem/db review with YZ
  • This week
    • IT infrastcuture outlook
    • ATT monthly
    • SWG reports
    • dms mongodb split
    • ks ims/dms f34

YZ

  • Previous:
    • OSCO/Cravat program upgrade (big performance boost/new functionality, new version ready).
    • Some support work after power outage on Tuesday.
    • MODS raritan KVM installation NOT yet completed (additional parts ordered).
    • MODS2 Linux machine lost a disk in its RAID1 array (OSU team will examine it later).
    • Some TMS data discussion with Andrew and Heejoo.
    • Some MODS WCS work and discussion with Olga and Dave T.
    • Some discussion of TCS refmem refactoring with MB.
  • Next:
    • Some OSCO work (rename the github repo, maybe rpm build and new release, and additional work/exploration if have time).
    • Work on DMS TCS refmem refactoring.
    • MODS KVM installation scheduled to complete.
    • Some misc. MODS work and discussions.
    • Re-start some TMS work if have time.
    • Safety training.

DE

  • Final hardware distribution save for Glenn's Ethernet adapter
  • CDW quotes
  • Trying to get MS Server software; looks like the org chart is in flux so we may be caught behind the shuffle. No phone number for contact and now he's directly under the CIO, so I'm not sure what the best next step is.
  • Autodesk licensing still broken
  • Smithwright software installs; no response so I assume it all went fine.
  • Prep for infrastructure expenditure meetings this week

PSG

  • Last week
    • Attempt to edit D425 firmware With Gustavo.
    • Scout software can only write to the controller, editing is not possible.
    • In order to modify the firmware a project file is needed, from which the updated firmware will be generated.
  • This week
    • Talk with Wolfgang to determine if this project file(s) exist.
    • Modify and use them to update the controller.

GCE

  • Last week
    • Set up work laptop
    • Benefits and Training
    • Ride-along with Dave Thompson during LUCI1 telescope checkout
    • Begin reviewing LUCI software handoff documentation
  • This week
    • Finish benefits and training
    • Discuss LUCI priorities with Dave Thompson
    • Discuss LUCI alert scripts with Jenny Power
    • Review and understand IIF
    • Get access to Development VMs and LUCI Simulator
    • Create LUCI simulator on local machine

BM

  • Last week
    • Onboarding, keys and building access, laptop, relocating
    • Started AO tutorial sessions with Al
  • This week
    • Back in Denver, will be in Tucson September 12
    • Begin reviewing DMS repository, wiki
    • Continue reading AO documentation (videos and presentations on wiki)
    • HSI training

SH

  • power down/power up
    • Weatherstation needed two reboots
    • Mount points on dms.mountain, tcs1, web
  • VPN RADIUS to Google SSL issue
  • Mountain calls (Thursday, Friday, Saturday morning).
  • LDAP

next:

  • IT planning meetings, writing
Topic revision: r8 - 23 Aug 2021, DavidEdmeades
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback